OpenStudy (anonymous):

i need help!!! Please If f(x)=2x^2-x-6 and g(x)= x^2-4, find f(x)/g(x) Answers: 2x+3/x-2 2x-3/x+2 2x+3/x+2 2x-3/x-2

OpenStudy (zpupster):

if you factor these one might cancel

OpenStudy (anonymous):

I dont get it @zpupster

OpenStudy (zpupster):

2x^2-x-6/x^2-4 here we see the denominator is the difference of 2 squares x^2-4= (x-2)(x+2) 2x^2-x-6 what does this factor into?

OpenStudy (anonymous):

umm 3x^2-6? im not good in math at al please dont laugh lol @zpupster

OpenStudy (zpupster):

step by step (2x (x

OpenStudy (zpupster):

(2x 3) (x 2) we need to find the signs

OpenStudy (anonymous):

but dont we have to pair them up @zpupster

OpenStudy (zpupster):

(2x -3) (x +2) = 2x^2 + x -6 that is not right so switch the signs (2x +3) (x -2) = 2x^2 - x -6 that is right

OpenStudy (zpupster):

(2x +3) (x -2)/(x+2)(x-2) what cancels???

OpenStudy (anonymous):

ohh okay i kinda get it now @zpupster

OpenStudy (anonymous):

one of the the (x-2) canaels right? @zpupster

OpenStudy (anonymous):

cancels*

OpenStudy (zpupster):

yes you can see your answer now. I ma sorry I had to leave yesterday.

OpenStudy (anonymous):

thank you its okay @zpupster